Circular No. 6521

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 22P/KOPFF G. Cremonese, Astronomical Observatory, Padua; and K. Rembor,

Max-Planck-Institut für Aeronomie, Lindau, report: “Using the 1.5-m Danish telescope at the European Southern Observatory on Dec. 6 we observed this comet to be split into at least three components. Our images reveal a main nucleus, showing strong jet activity, and a complex of fainter fragments toward p.a. 70^{o} that also show activity; there is also an elongated coma toward p.a. 200^{o}. In a Bessel R image the separation between the main nucleus and the fragments was ∼ 40′′ (14 000 km); a Bessel B image obtained 7 min later shows quite a different appearance, the separation ∼ 60′′ (20 000 km). The fragments did not appear in observations on the three following nights with the 2.2-m telescope, and the comet then showed a strong fan close to the sunward direction. On B, V and R images obtained with the 3.6-m telescope on Nov. 13 and 14 the comet showed strong dust activity (much more so than in December). By applying the ring-masking method it becomes evident that this activity was characterized by a fan toward p.a. 180^{o} and a prominent asymmetry in the inner coma.”

Ephemeris from the orbital elements on MPC 27080:

1996/97 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 13 22^{h}41.^{m}50 −12^{o}43.^{′}6 2.235 2.211 75.^{o}8 25.^{o}6 15.2
23 22 58.66 −10 54.7 2.405 2.270 70.2 24.1 15.7
Jan. 2 23 15.60 − 9 04.1 2.575 2.330 64.6 22.4 16.1
12 23 32.31 − 7 12.9 2.742 2.389 59.0 20.7 16.5

COMET C/1996 Q1 (TABUR) The Nov. 23 observation on IAUC 6513 has not been confirmed by

others, and no satisfactory astrometry has been possible since Oct. 16, i.e., before the comet rapidly faded and dramatically dispersed as noted on IAUC 6499. If this is indeed the secondary component of a comet that split 2900 years ago to be nongravitationally separated now from the primary by more than eight years (cf. IAUC 6464), such behavior may not be particularly surprising. Ephemeris continuation (orbit on MPC 28052):

1996/97 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 13 16^{h}03.^{m}50 +18^{o}07.^{′}5 1.545 1.107 45.^{o}5 39.^{o}4 12.4
23 16 09.45 +15 18.8 1.638 1.225 48.2 36.7 13.0
Jan. 2 16 13.81 +13 02.4 1.700 1.352 52.6 35.3 13.5
12 16 16.41 +11 13.1 1.732 1.482 58.7 34.5 13.9
1996 December 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden